May 4, 2017 Calorimeter

From GlueXWiki
Jump to: navigation, search

Video Conferencing Information

Meeting Time: 11:00 a.m.

  1. To join via a Web Browser, go to the page [1]
  2. To join via Polycom room system go to the IP Address: ( and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone in CC L207 is 757-269-7084 (usual room)
  • JLab Phone: in CC F326 is 757-269-6460
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Action Items

  1. Disable red LED string in dark room to avoid leaving it on in the future. (Adesh)


  1. FCAL HDFCAL log book
  2. BCAL HDBCAL log book

Tentative Agenda

  1. Announcements
  2. Collaboration Meeting May 15-17, 2017
  3. Action Items
  4. Activities during the down
  5. FCAL update
  6. BCAL update
  7. Calibrations
  8. Reconstruction and Simulation
  9. NIM articles
  10. Any other business


Attendees: Elton, Adesh, Mark (JLab); Will (CMU); Sean (NW); Christina (Athens); Matt (IU); Zisis (Regina)

  1. Announcements
  2. Collaboration Meeting May 15-17, 2017
  3. Action Items
    • Matt: Add maintenance of FCAL bases to action item
  4. Activities during the down
    • Elton: We have had discussions with Chris Stanislav regarding repair of some BCAL channels.
    • Zisis: Tegan will arrive on May 17 and Zisis on May 23 and would like to take more LED data, especially turning off selected HV for better understanding of individual channels.
  5. FCAL update (Adesh)
    1. 4 ns shifts
      • Offsets have been determined relative to a reference run. At first he found 0, +/- 4 ns offsets. Speaking to Mark, we only expect 0 and 4 ns offsets. Adjustment of the reference run allows al offsets to be set to 0, 4 ns. This is being implemented.
      • A second pass will determine offsets of one channel relative the others by referencing these to the RF signal.
      • These time offsets are expected to be completed in about 2 weeks, likely not before the production launch.
      • Mark: Will need a new table in CCDB to hold the offsets
      • Mark: What is the timing used for? Matt: hit selection based on timing, e.g. eliminating EM showers that are out of time. Sean: Timing selection are being made to the RF bucket. Matt: Uncomfortable making such tight timing cuts on the FCAL at this stage of calibration.
    2. Low intensity runs / normalization
      • Finds that for the low intensity runs, the pi0 mass peak shifts down by 4-5 MeV. This is not understood, but is confirmed by LED data, which shows a 5% drop in intensity. There also appears a non-trial geometry dependence to this systematic. During that period the pedestals and HV remain constant.
    3. time walk corrections
      • The time-walk corrections to correct for slewing in the FCAL timing has been implemented in Adesh's local branch. Source of slewing is unclear but corrected away.
  6. BCAL update (Mark)
    1. 4 ns offsets
      • Offsets for 4 ns have been completed and entered into the CCDB run-by-run
    2. Higher level calibrations (attenuation length, gain ratio, position offsets, time difference offsets and time sum offsets) completed
      • These calibrations have been extracted once and seem to be valid over the entire run
      • Sean initially saw some discrepancies in PID variables, but they seem to be related to his software installation
      • All of Mark's calibrations are being used in the current monitoring launches.
    3. LED timing (Elton)
      • Timing offsets using the time difference between (down-up) were derived for runs available on /cache. They show good agreement with Mark's constants except for one run.
      • All 4ns offsets have been applied to the downstream end, as implied by comparisons to TDCs as well as raw times compared to the trigger.
      • A selection cut in the DBCALPoint_factory was eliminating hits in layers 4 and 3. It has been removed from the point factory. Will verified that it does not have any consequences on shower reconstruction as equivalent cuts are implemented there.
      • Elton: The ambiguities created by the 4 ns offsets would greatly simplify calibrations if they could be fixed in the hardware.
    4. List of problem channels (Elton, Mark and Ahmed)
    5. Sean: Encourages us to writeup the calibration procedures when there is time.
  7. Calibrations
    1. Gains BCAL (Will)
      • Are there any calibrations that still need uploading to the data base? Mark: No.
      • Gain constants need revision after all the recent changes to calibrations and constants
      • Gain constants reflect changes in the gain ratios of about 15%, leading to changes in the pi0 mass by about 5 MeV
      • Iterations are converging and expect to have new gain constants by end of Friday
      • Are the constants valid for all runs or just the low intensity? Mark: They should be valid for both low and high intensity periods.
      • Elton: Are you using the matrix inversion method or just adjustment of pi0 peak? Will: Using pi0 peak adjustment, which worked better for this period
      • Tegan has been making progress on calibration with an exclusive pi0 sample.
      • Layer 4 calibrations have been taken from Andrei's cosmic-ray calibration after adjusting for an overall normalization. No direct effect is seen in the pi0 peak as most showers do not spill into layer 4. However, these are probably a better set of constants for this layer.
    2. Monitoring (Ahmed)
      • Ran into some merging issues with BCAL_LED plugin. This was resolved by changing some histograms to root trees. Since these are mostly diagnostic tools, it should not affect the monitoring process.
      • This plugin will no longer be run as part of the launch. It will be run over skimmed files, so changes do not need to be fully compatible with other plugins during the launch.
      • Elton: when are all the skims going to be run? Sean: They have been completed and located in /cache/halld/RunPeriod-2017-01/calib/ver01/BCAL-LED but they may have become unpinned.
  8. Reconstruction and Simulation (Sean)
    • Need to start efforts to tune geant4 . BCAL seems to be in good shape. He has been in touch with Will, who has made adjustments of the MC to data for geant3. He will send files to Will to check the new simulation.
    • Time shift in HDGeant4 FCAL showers
    • Finds timing shift of about 1-2 ns in the FCAL shower times. It is probably related to the depth correction. This needs to be investigated.
    • Elton: Have the momentum resolutions been resolved for geant4? Sean: No, but at the moment addressing easier problems first. Tracking will come later.
  9. NIM articles
    • Zisis: BCAL NIM is progressing by producing figures from the Regina work that could be used in a publication. Next will need sections from Will and Mark.
  10. Any other business